OpenStudy (anonymous):

I am having issues understanding the following question Find the difference between 2 and the squares of ... (see reply) Could someone explain how one would start working on a question like this?

OpenStudy (anonymous):

\[\frac{ 1 }{ 1 },\frac{ 3 }{ 2 }, \frac{ 7 }{ 5 }, \frac{ 17 }{ 12 }, \frac{ 41 }{ 29 }, \frac{ 99 }{ 70 }\] a) Verify that successive terms of the sequence stand in relation to each other as m/n does to (m + 2n)/(m+n), b) If m/n is a good approximation of \[\sqrt{2}\] then (m + 2n)/(m+n) is a better one, and that the errors in the two cases are in opposite directions. c) Find the next three terms of the above sequence (cit: Glyn James, Modern Engineering Mathematics, exr. 1.2.3.5a-c)

OpenStudy (anonymous):

Wrong Area. Sorry. :|